Gunns Inc. issues 15,000 shares of $1 par value common stock and 25 shares of $1,000...

  1. Gunns Inc. issues 15,000 shares of $1 par value common stock and 25 shares of $1,000 par value, 6% preferred stock to a private investor for $630,000.  The fair value of the common stock is $40 per share and the fair value of the preferred stock is $1200 per share.
  1. Prepare the journal entry to record the transaction assuming that the fair market values (FMV) for both the common and preferred stock are known and shown below.
Answer #1
Journal entry in the books of Gunns Inc.
Dr/Cr Particulars Debit $ Credit $
Debit Cash/Bank       630,000
Credit Common Stock         15,000
Credit 6% Preferred Stock         25,000
Credit Paid in excess of par- Common Stock (Note:-1)       585,000
Credit Paid in excess of par- 6% Preferred Stock (Note:-2)           5,000
(To record the issue of common stock face value $1 and 6% preferred stock $ 1000 per share at fair market value to a investor)
Note:-1
Cell Reference Particulars Details
A No of common stock                                                                             15,000
B Face Value of common stock                                                                                        1
C Marker value of common stock                                                                                     40
D=C-B Paid in excess of par- Common Stock per share                                                                                     39
E=D*A Paid in excess of par- Common Stock                                                                           585,000
Note:-2
Cell Reference Particulars Details
A No of 6% Preferred stock                                                                                     25
B Face Value of 6% Preferred stock                                                                               1,000
C Marker value of 6% Preferred stock                                                                               1,200
D=C-B Paid in excess of par- 6% Preferred stock per share                                                                                   200
E=D*A Paid in excess of par- 6% Preferred stock                                                                               5,000